# CASE: Contrastive Activation for Saliency Estimation

### v2

@@ Abstract
I Introduction
As deep learning systems become more widely deployed, explanation methods are increasingly relied upon to justify, debug, and evaluate model predictions. Activation-based approaches such as Class Activation Mapping (CAM) [1] and Gradient-Weighted Class Activation Mapping (Grad-CAM) [2] are among the most widely used tools for interpreting convolutional neural networks (CNNs). These methods aim to highlight features in the input that are most relevant to a particular class prediction. However, despite their popularity, such saliency methods often exhibit a critical failure mode: they produce nearly identical explanations for different class labels on the same input. [3, 4]. This redundancy undermines their utility in settings where distinguishing between predictions is essential, such as auditing errors, understanding misclassifications, or building trust in model behavior. [5, 6, 7, 8, 9]
This concern is echoed in human-centered studies such as HIVE [10], which show that visual explanations can increase user trust even when the model is wrong, raising the risk that visually plausible but semantically empty saliency maps may mislead stakeholders. They also observe that existing methods often produce indistinguishable explanations for correct and incorrect predictions, suggesting a broader insensitivity to class label. In high-stakes applications, such as diagnosis or safety monitoring, this insensitivity directly compromises the value of explanations as a safeguard against error. Notably, real-world failures of saliency methods have been documented in medical imaging, where models mistakenly rely on dataset artifacts, such as text tokens or positional cues, rather than pathology features, leading to clinically misleading saliency maps. [9].
@@ I Introduction
Class-Specificity: Explanations should reflect features that are distinctive to the target class. Saliency maps for semantically disjoint labels (e.g., airplane vs. barbell) should differ more than those for related ones (e.g., terrier vs. poodle). Bau et al. [11] demonstrate that CNN units can align with human-interpretable concepts, suggesting that class-specific representations are both learnable and quantifiable.
@@
II Related Work
II-A Contrastive Explanations
A number of prior methods have explored contrastive reasoning to enhance interpretability. CWOX [16] formulates explanation as a clustering problem over confusion sets, requiring class-specific explanations to differ from those of their nearest neighbors in a latent tree. While conceptually appealing, CWOX relies on HLTA-derived cluster structures and explicit contrast baselines, which may introduce brittleness or restrict applicability. [17]
Contrastive Layer-wise Relevance Propagation (C-LRP) [18] extends the LRP [19] framework to compare class-specific relevance maps to those of reference classes. However, C-LRP remains tied to backpropagation-based relevance signals and does not extend to activation-based methods like CAM.
@@ II-A Contrastive Explanations
II-B Saliency Methods and Class Insensitivity
Grad-CAM [2] and its variants (e.g., Grad-CAM++ [20], AblationCAM [21], ScoreCAM [22], LayerCAM [23]) have become standard tools for visualizing model predictions. These methods produce spatial heatmaps based on either gradient or activation weighting. However, several studies have highlighted their limitations with respect to faithfulness and class sensitivity.
Adebayo et al. [24] demonstrate that certain saliency methods are invariant to randomization of model weights or labels, failing what they term ”sanity checks.”. Their work introduces the notion of model-insensitivity, where attribution maps remain unchanged even as the model parameters are randomized. While this critique reveals concerning invariance patterns, it targets variation across models, not across class predictions within the same model.
@@ II-B Saliency Methods and Class Insensitivity
II-C Discriminative Features and Exclusivity
Beyond attribution intensity, several works have investigated the exclusivity of features with respect to class labels. Kalibhat et al. [25] define discriminative features as those with positive gradient contributions toward a class score:
@@ II-C Discriminative Features and Exclusivity
II-D Agreement Metrics and Evaluation of Saliency
Several metrics have been proposed to evaluate saliency maps based on consistency and selectivity. Krishna et al. [26] introduce the notion of the disagreement problem, observing that saliency maps for different class labels often highlight overlapping regions. They propose overlap-based metrics to assess class-separability in explanations. We adopt and extend this framework by applying the Wilcoxon signed-rank test to determine whether saliency methods systematically fail to distinguish between top-1 and top-2 predictions. This offers a statistical lens on class sensitivity that complements visual inspection.
II-E Summary
In summary, while prior work has surfaced critical limitations of saliency methods, such as insensitivity to model or data [24], or failure to isolate class-specific features [26, 25], few methods provide both class-distinctiveness and architectural compatibility. Our work addresses this gap by quantifying class overlap in existing methods and introducing a contrastive refinement tailored to activation-based explanations. This yields clearer, more class-sensitive visualizations and aligns explanatory focus with the model’s internal decision boundaries.
III Contrastive Activation for Saliency Estimation (CASE)
III-A Overview
@@
III-B Contrastive Gradients
Prior work has observed that many saliency methods produce overlapping explanations across distinct class labels [26], limiting their ability to convey what distinguishes one class from another. In CASE, we begin by computing the gradient of the class score with respect to the activation map for the target class . We then suppress the components of this gradient that are aligned with an average gradient vector derived from a contrast set, i.e., the classes most frequently confused with . This orthogonalization removes shared attribution directions and emphasizes features that are uniquely predictive of .
III-C Confusion-Based Contrast Set Selection
@@
III-D Uniquely Discriminative Features
@@ IV-B RQ1: Class Sensitivity
To quantify this behavior, we use the feature agreement metric proposed by [26], which computes the proportion of shared features among the top- salient pixels:
@@ IV-C RQ2: Explanation Faithfulness
IV-D Models, Datasets, and Methods
Our evaluation includes four pre-trained image classification models: ResNet-50, VGG-19, DenseNet-201, and ConvNeXt-Large [27, 28, 29]. Models are evaluated on two standard datasets: ImageNet [30], which offers broad category diversity across 1,000 classes, and CIFAR-100 [31], which contains 100 object categories drawn from 20 coarse-grained superclasses. CIFAR-100 consists of low-resolution natural images () covering a wide range of everyday concepts such as animals, vehicles, and household objects.
@@ IV-D Models, Datasets, and Methods
V Results
V-A Class Sensitivity Analysis
@@
Table I reports the -values. For each entry, if , we reject the null hypothesis and conclude that the method produces significantly different saliency maps for the top-1 and top-2 class labels. If , we fail to reject the null and cannot conclude that the explanations differ.
@@
V-A1 Robustness to Training Seeds (DenseNet on CIFAR-100)
@@
V-B Ablation Study: Channel Sparsity and Discriminative Structure
@@
Notably, when we re-ran RQ1 using features.norm5, the final normalization layer in DenseNet-201, instead of the final convolutional layer, the results were markedly different. As shown in Table II, most saliency methods failed to reject the null hypothesis, indicating that their explanations did not significantly differ across class labels. This stands in sharp contrast to the results reported in Table I for the convolutional layer, where nearly all methods achieved significant class sensitivity. These findings reinforce the conclusion that saliency quality is highly sensitive to layer choice, and that DenseNet’s class-distinct explanations are driven by the sparse and selective representations encoded in its final convolutional layer. Notably, CASE retains its class sensitivity even under this more diffuse, high-activation regime.
@@ V-C Explanation Fidelity
Table III presents the results across all four architectures. On ConvNeXt, CASE significantly outperforms AblationCAM and LayerCAM, while showing no statistical difference from Grad-CAM, Grad-CAM++, or ScoreCAM. On DenseNet and ResNet, no baseline method differs significantly from CASE, indicating comparable levels of confidence drop. On VGG, three methods, Grad-CAM++, LayerCAM, and ScoreCAM, yield significantly higher drops than CASE, while the remaining methods are not statistically distinguishable.
@@
VI Discussion
VI-A RQ1: Class Sensitivity of Saliency Methods
The results in Table I demonstrate that the ability of saliency methods to produce class-specific explanations varies substantially across model architectures. DenseNet stands out for its consistent rejection of the null hypothesis across all tested saliency methods, indicating robust class-distinct attribution. In contrast, architectures like VGG and ResNet fail to produce class-sensitive maps under several methods, notably Grad-CAM and AblationCAM. ScoreCAM achieves broader consistency, performing well on both DenseNet and ResNet, while methods such as LayerCAM and Grad-CAM++ show stronger dependency on the underlying architecture.
@@ VI-A RQ1: Class Sensitivity of Saliency Methods
The ablation in Table II underscores that effective attribution depends not only on proximity to the classifier, but also on the structural selectivity of the chosen layer. These results motivate the need to treat attribution layer selection as a principled design decision, rather than a fixed convention.
@@
VI-B RQ2: Explanation Fidelity of Saliency Methods
The results in Table III indicate that CASE achieves statistically significant confidence drops compared to AblationCAM and LayerCAM on ConvNeXt, and performs comparably to the strongest baselines on DenseNet and ResNet. On VGG, several methods: Grad-CAM++, LayerCAM, and ScoreCAM, achieve higher confidence drops than CASE, though only Grad-CAM++ and LayerCAM are significant.
@@ VI-B RQ2: Explanation Fidelity of Saliency Methods
VII Conclusion and Future Directions
@@ VIII Limitations
Preprint Notice
This work is currently under review at IEEE Transactions on Neural Networks and Learning Systems (TNNLS).
@@ Appendix A Appendix
A-A Agreement Score Distributions by Method
To complement the statistical analysis reported in Section V-A, we visualize the distribution of class agreement scores for each saliency method across all tested architectures. These histograms represent the percentage overlap between top-1 and top-2 saliency maps, aggregated over all validation images.
@@ A-A Agreement Score Distributions by Method
We present these distributions for each of the four architectures evaluated in the main paper: VGG, ResNet, ConvNeXt, and DenseNet. For DenseNet, we include two variants: one using the final convolutional layer (conv2), and another using the final normalization layer (norm5). This dual analysis allows us to visualize how attribution quality shifts with layer selection, supporting our ablation findings in Section V-B.
VGG-19.
VGG serves as a strong baseline for class sensitivity analysis due to its simple, sequential architecture. As the attribution layer activates broadly across many channels, we expect higher top-1/top-2 agreement. The distributions below show that most methods yield class-insensitive explanations, with significant overlap above the 50% threshold.
@@
ResNet introduces residual connections that promote feature reuse and broader spatial coverage. These plots reveal that many saliency methods still produce class-indistinct maps, with distributions often centered near or above the 50% threshold. CASE, however, consistently shifts agreement lower, reflecting stronger separation of class-relevant features.
@@
ConvNeXt offers a modern convolutional backbone with enhanced channel mixing and normalization. Despite its design, several methods exhibit high agreement scores across class predictions, indicating limited class specificity. CASE maintains a clear advantage, with distributions skewed leftward below the threshold.
@@
This is the final convolutional layer used in our main experiments. As detailed in Section V-B, it is extremely sparse yet highly class-discriminative. The agreement distributions here confirm that most methods produce low-overlap explanations, consistent with the high class sensitivity reported in RQ1.
@@
This final normalization layer aggregates upstream activations and is located immediately before the classifier. As discussed in our ablation study, this layer activates nearly all channels uniformly. The agreement distributions below reveal that most methods fail to produce class-distinct saliency maps in this setting, highlighting the importance of attribution layer selection.

### v3

@@ Abstract
1 Introduction
As deep learning systems become more widely deployed, explanation methods are increasingly relied upon to justify, debug, and evaluate model predictions. Activation-based approaches such as Class Activation Mapping (CAM) (Zhou et al., 2015) and Gradient-Weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2016) are among the most widely used tools for interpreting convolutional neural networks (CNNs). These methods aim to highlight features in the input that are most relevant to a particular class prediction. However, despite their popularity, such saliency methods often exhibit a critical failure mode: they produce nearly identical explanations for different class labels on the same input. (Hooker et al., 2019; Jalwana et al., 2021). This redundancy undermines their utility in settings where distinguishing between predictions is essential, such as auditing errors, understanding misclassifications, or building trust in model behavior. (Alqaraawi et al., 2020; Lakkaraju et al., 2019; Lage et al., 2019; Mohseni et al., 2021; Tjoa and Guan, 2021)
This concern is echoed in human-centered studies such as HIVE (Kim et al., 2021), which show that visual explanations can increase user trust even when the model is wrong, raising the risk that visually plausible but semantically empty saliency maps may mislead stakeholders. They also observe that existing methods often produce indistinguishable explanations for correct and incorrect predictions, suggesting a broader insensitivity to class label. In high-stakes applications, such as diagnosis or safety monitoring, this insensitivity directly compromises the value of explanations as a safeguard against error. Notably, real-world failures of saliency methods have been documented in medical imaging, where models mistakenly rely on dataset artifacts, such as text tokens or positional cues, rather than pathology features, leading to clinically misleading saliency maps. (Tjoa and Guan, 2021).
@@ 1 Introduction
Class-Specificity: Explanations should reflect features that are distinctive to the target class. Saliency maps for semantically disjoint labels (e.g., airplane vs. barbell) should differ more than those for related ones (e.g., terrier vs. poodle). Bau et al. (2017) demonstrate that CNN units can align with human-interpretable concepts, suggesting that class-specific representations are both learnable and quantifiable.
@@
2 Related Work
2.1 Contrastive Explanations
A number of prior methods have explored contrastive reasoning to enhance interpretability. CWOX (Xie et al., 2023) formulates explanation as a clustering problem over confusion sets, requiring class-specific explanations to differ from those of their nearest neighbors in a latent tree. While conceptually appealing, CWOX relies on HLTA-derived cluster structures and explicit contrast baselines, which may introduce brittleness or restrict applicability. (Chen et al., 2016)
Contrastive Layer-wise Relevance Propagation (C-LRP) (Gu et al., 2018) extends the LRP (Bach et al., 2015) framework to compare class-specific relevance maps to those of reference classes. However, C-LRP remains tied to backpropagation-based relevance signals and does not extend to activation-based methods like CAM.
@@ 2.1 Contrastive Explanations
2.2 Saliency Methods and Class Insensitivity
Grad-CAM (Selvaraju et al., 2016) and its variants (e.g., Grad-CAM++ (Chattopadhay et al., 2018), AblationCAM (Desai and Ramaswamy, 2020), ScoreCAM (Wang et al., 2019), LayerCAM (Jiang et al., 2021)) have become standard tools for visualizing model predictions. These methods produce spatial heatmaps based on either gradient or activation weighting. However, several studies have highlighted their limitations with respect to faithfulness and class sensitivity.
Adebayo et al. (2018) demonstrate that certain saliency methods are invariant to randomization of model weights or labels, failing what they term ”sanity checks.”. Their work introduces the notion of model-insensitivity, where attribution maps remain unchanged even as the model parameters are randomized. While this critique reveals concerning invariance patterns, it targets variation across models, not across class predictions within the same model.
@@ 2.2 Saliency Methods and Class Insensitivity
2.3 Discriminative Features and Exclusivity
Beyond attribution intensity, several works have investigated the exclusivity of features with respect to class labels. Kalibhat et al. (2023) define discriminative features as those with positive gradient contributions toward a class score:
@@ 2.3 Discriminative Features and Exclusivity
2.4 Agreement Metrics and Evaluation of Saliency
Several metrics have been proposed to evaluate saliency maps based on consistency and selectivity. Krishna et al. (Krishna et al., 2022) introduce the notion of the disagreement problem, observing that saliency maps for different class labels often highlight overlapping regions. They propose overlap-based metrics to assess class-separability in explanations. We adopt and extend this framework by applying the Wilcoxon signed-rank test to determine whether saliency methods systematically fail to distinguish between top-1 and top-2 predictions. This offers a statistical lens on class sensitivity that complements visual inspection.
2.5 Summary
In summary, while prior work has surfaced critical limitations of saliency methods, such as insensitivity to model or data (Adebayo et al., 2018), or failure to isolate class-specific features (Krishna et al., 2022; Kalibhat et al., 2023), few methods provide both class-distinctiveness and architectural compatibility. Our work addresses this gap by quantifying class overlap in existing methods and introducing a contrastive refinement tailored to activation-based explanations. This yields clearer, more class-sensitive visualizations and aligns explanatory focus with the model’s internal decision boundaries.
3 Contrastive Activation for Saliency Estimation (CASE)
3.1 Overview
@@
3.2 Contrastive Gradients
Prior work has observed that many saliency methods produce overlapping explanations across distinct class labels (Krishna et al., 2022), limiting their ability to convey what distinguishes one class from another. In CASE, we begin by computing the gradient of the class score with respect to the activation map for the target class . We then suppress the components of this gradient that are aligned with an average gradient vector derived from a contrast set, i.e., the classes most frequently confused with . This orthogonalization removes shared attribution directions and emphasizes features that are uniquely predictive of .
3.3 Confusion-Based Contrast Set Selection
@@
3.4 Uniquely Discriminative Features
@@ 4.2 RQ1: Class Sensitivity
To quantify this behavior, we use the feature agreement metric proposed by (Krishna et al., 2022), which computes the proportion of shared features among the top- salient pixels:
@@ 4.3 RQ2: Explanation Faithfulness
4.4 Models, Datasets, and Methods
Our evaluation includes four pre-trained image classification models: ResNet-50, VGG-19, DenseNet-201, and ConvNeXt-Large (He et al., 2015; Liu et al., 2022; Huang et al., 2016). Models are evaluated on two standard datasets: ImageNet (Deng et al., 2009), which offers broad category diversity across 1,000 classes, and CIFAR-100 (Krizhevsky, 2009), which contains 100 object categories drawn from 20 coarse-grained superclasses. CIFAR-100 consists of low-resolution natural images () covering a wide range of everyday concepts such as animals, vehicles, and household objects.
@@ 4.4 Models, Datasets, and Methods
5 Results
5.1 Class Sensitivity Analysis
@@
Table 1 reports the -values. For each entry, if , we reject the null hypothesis and conclude that the method produces significantly different saliency maps for the top-1 and top-2 class labels. If , we fail to reject the null and cannot conclude that the explanations differ.
@@
5.1.1 Robustness to Training Seeds (DenseNet on CIFAR-100)
@@
5.2 Ablation Study: Channel Sparsity and Discriminative Structure
@@
Notably, when we re-ran RQ1 using features.norm5, the final normalization layer in DenseNet-201, instead of the final convolutional layer, the results were markedly different. As shown in Table 2, most saliency methods failed to reject the null hypothesis, indicating that their explanations did not significantly differ across class labels. This stands in sharp contrast to the results reported in Table 1 for the convolutional layer, where nearly all methods achieved significant class sensitivity. These findings reinforce the conclusion that saliency quality is highly sensitive to layer choice, and that DenseNet’s class-distinct explanations are driven by the sparse and selective representations encoded in its final convolutional layer. Notably, CASE retains its class sensitivity even under this more diffuse, high-activation regime.
@@ 5.3 Explanation Fidelity
Table 3 presents the results across all four architectures. On ConvNeXt, CASE significantly outperforms AblationCAM and LayerCAM, while showing no statistical difference from Grad-CAM, Grad-CAM++, or ScoreCAM. On DenseNet and ResNet, no baseline method differs significantly from CASE, indicating comparable levels of confidence drop. On VGG, three methods, Grad-CAM++, LayerCAM, and ScoreCAM, yield significantly higher drops than CASE, while the remaining methods are not statistically distinguishable.
@@
6 Discussion
6.1 RQ1: Class Sensitivity of Saliency Methods
The results in Table 1 demonstrate that the ability of saliency methods to produce class-specific explanations varies substantially across model architectures. DenseNet stands out for its consistent rejection of the null hypothesis across all tested saliency methods, indicating robust class-distinct attribution. In contrast, architectures like VGG and ResNet fail to produce class-sensitive maps under several methods, notably Grad-CAM and AblationCAM. ScoreCAM achieves broader consistency, performing well on both DenseNet and ResNet, while methods such as LayerCAM and Grad-CAM++ show stronger dependency on the underlying architecture.
@@ 6.1 RQ1: Class Sensitivity of Saliency Methods
The ablation in Table 2 underscores that effective attribution depends not only on proximity to the classifier, but also on the structural selectivity of the chosen layer. These results motivate the need to treat attribution layer selection as a principled design decision, rather than a fixed convention.
@@
6.2 RQ2: Explanation Fidelity of Saliency Methods
The results in Table 3 indicate that CASE achieves statistically significant confidence drops compared to AblationCAM and LayerCAM on ConvNeXt, and performs comparably to the strongest baselines on DenseNet and ResNet. On VGG, several methods: Grad-CAM++, LayerCAM, and ScoreCAM, achieve higher confidence drops than CASE, though only Grad-CAM++ and LayerCAM are significant.
@@ 6.2 RQ2: Explanation Fidelity of Saliency Methods
7 Conclusion and Future Directions
@@ Appendix A Appendix
A.1 Agreement Score Distributions by Method
To complement the statistical analysis reported in Section 5.1, we visualize the distribution of class agreement scores for each saliency method across all tested architectures. These histograms represent the percentage overlap between top-1 and top-2 saliency maps, aggregated over all validation images.
@@ A.1 Agreement Score Distributions by Method
We present these distributions for each of the four architectures evaluated in the main paper: VGG, ResNet, ConvNeXt, and DenseNet. For DenseNet, we include two variants: one using the final convolutional layer (conv2), and another using the final normalization layer (norm5). This dual analysis allows us to visualize how attribution quality shifts with layer selection, supporting our ablation findings in Section 5.2.
VGG.
serves as a strong baseline for class sensitivity analysis due to its simple, sequential architecture. As the attribution layer activates broadly across many channels, we expect higher top-1/top-2 agreement. The distributions in Figure 7 show that most methods yield class-insensitive explanations, with significant overlap above the 50% threshold.
@@
ResNet introduces residual connections that promote feature reuse and broader spatial coverage. ( Figure 8) reveals that many saliency methods still produce class-indistinct maps, with distributions often centered near or above the 50% threshold. CASE, however, consistently shifts agreement lower, reflecting stronger separation of class-relevant features.
@@
ConvNeXt offers a modern convolutional backbone with enhanced channel mixing and normalization. Despite its design, several methods exhibit high agreement scores across class predictions (see Figure 9), indicating limited class specificity. CASE maintains a clear advantage, with distributions skewed leftward below the threshold.
@@
This is the final convolutional layer used in our main experiments. As detailed in Section 5.2, it is extremely sparse yet highly class-discriminative. The agreement distributions here (see Figure 10) confirm that most methods produce low-overlap explanations, consistent with the high class sensitivity reported in RQ1.
@@
This final normalization layer aggregates upstream activations and is located immediately before the classifier. As discussed in our ablation study, this layer activates nearly all channels uniformly. The agreement distributions ( Figure 11) reveal that most methods fail to produce class-distinct saliency maps in this setting, highlighting the importance of attribution layer selection.